# Latent Concept-based Explanation of NLP Models

### v1

@@ 2.4 PlausiFyer
3 Experimental Setup
Data
We use Parts-of-Speech (POS) Tagging and Sentiment Classification (Sentiment) tasks for our experiments. The former is a sequence labeling task, where every word in the input sentence is assigned a POS tag, while the latter classifies sentences into two classes representing Positive and Negative sentiment. We use the Penn TreeBank dataset Marcus et al. (1993) for POS and the ERASER Movie Reviews dataset Pang and Lee (2004); Zaidan and Eisner (2008) for Sentiment. The POS dataset consists of 36k, 1.8k and 1.9k splits for train, dev and test respectively and 44 classes.
The ERASER movie review dataset consists of labeled paragraphs with human annotations of the words and phrases. We filter sentences that have a word/phrase labeled with sentiment and create a sentence-level sentiment classification dataset. The dataset consists of 9.4k positive and 8.6k negative instances. We randomly split it into 13k, 1.5k and 2.7k splits for train, dev and test respectively.
Models
We fine-tune 12-layered pre-trained models; BERT-base-cased Devlin et al. (2019), RoBERTa-base Liu et al. (2019) and XLM-Roberta Conneau et al. (2020) using the training datasets of the tasks. We use transformers Wolf et al. (2020) with the default settings and hyperparameters. Task-wise performance of the models is provided in Appendix Tables 4 and  5.
Module-specific hyperparameters
When extracting the activation and/or attribution of a word, we average the respective value over the word’s subword units. We optimize the number of clusters $K$ for each dataset as suggested by Dalvi et al. (2022). We use $K=600$ (POS) and $K=400$ (Sentiment) for ConceptDiscoverer.
4 Evaluation
We perform a qualitative evaluation, a human evaluation and a module-level evaluation of LACOAT to measure its correctness and efficacy.
4.1 Qualitative Evaluation
4.1.1 Evolution of Concepts
LACOAT generates the explanation for each layer with respect to a prediction. The layer-wise explanation shows the evolution of concepts in making the prediction. Figure 2 shows layers 0, 6 and 12’s latent concept of the most attributed input token for RoBERTa fine-tuned on the sentiment task (see App Fig 5 for other examples). We found that the initial layer latent concepts do not always align with the sentiment of the input instance and may represent a general language concept. For instance, Figure 2(a) shows the concept of comparative and superlative adjectives of both positive and negative sentiments and is not limited to representing the negative sentiment of the most attributed word. In the middle layers, the latent concepts evolved into concepts that align better with the sentiment of the input sentence as can be seen in Figure 2(b). For instance, the latent concept of Figure 2(b) shows a mix of adjectives and adverbs of negative sentiment, i.e. aligned with the sentiment of the input sentence. In the sentiment task, the most attributed word in the last layer is [CLS] which resulted in latent concepts consisting of [CLS] representations of the most related sentences to the input. In such cases, we randomly pick five [CLS] instances from the latent concept and show their corresponding sentences in the figure (see Figure 2(c)). We found that the last layer’s latent concepts are best aligned with the input instance and its prediction and are the most informative explanation of the prediction. In the rest of the paper, we focus our analysis on the explanations generated using the last layer only and perform a human evaluation to evaluate their efficacy and correctness.
4.1.2 Analyzing Last Layer Explanations
Correct prediction with correct gold label
Wrong prediction with correct gold label
To study how the explanation would change if it is a correct prediction, we employ TextAttack Morris et al. (2020) to create an adversarial example of the sentence in Figure 3(b) that flips its prediction. The new sentence replaces “laughing” with “kidding” which has a similar meaning but flipped the prediction to a correct prediction. Figure 6 in the appendix shows the full explanation of the augmented sentence. With the correct prediction, the latent concept changed and the explanation clearly expresses a negative sentiment “… all express negative opinions and criticisms …" compared to the explanation of the wrongly predicted sentence.
Cross model analysis
LACOAT provides an opportunity to compare various models in terms of how they learned and structured the knowledge of a task. Figure 4 compares the explanation of RoBERTa (top) and XLMR (bottom) for identical inputs. Both models predicted the correct label. However, their latent concept based explanation is substantially different. RoBERTa’s explanation shows a large and diverse concept where many words are related to finance and economics. The XLMR’s latent concept is rather a small focused concept where the majority of tokens are units of measurement. It is worth noting that both models are fine-tuned on identical data.
Figure 4: Comparing explanation of RoBERTa (top) and XLMR (bottom)
4.1.3 Comparison to Input Feature based Explanation
In Figure 2, layer 0 serves as a direct comparison to the explanation provided by the attribution methods. Here, IG highlights the input word “worst" in the given example. While it informs the user of the most salient input token that effects the prediction, it does not provide insights into the particular facet the model has used in that particular example. Even if we compare the explanation of other layers like layer 6, the latent concept of the attributed word is more informative than the attributed word alone.
4.2 Human Evaluation
To evaluate the explanation generated by LACOAT empirically, we conduct a human evaluation using four annotators across 50 test samples. Specifically, given an explanation (e.g. Figure 3), three annotators are asked to answer the following five questions:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    Is the explanation relevant to the task at hand? (Yes/No)
Q1 evaluates whether LACOAT attributes the correct concept to a given prediction, while Q2 and Q3 measure the efficacy of LACOAT’s output in helping a user understand the prediction. Q4 and Q5 evaluates the output of PlausiFyer. They specifically separate out the cases where the explanation was accurate but irrelevant to the task at hand.
Labels Correct Samples Incorrect Samples All Samples
Annotation Fleiss $\kappa$
Q1 Yes/No 28 / 0 20 / 2 48 / 2 0.35
Q2 Helps/Neutral/Hinders 27 / 1 / 0 17 / 5 / 0 44 / 6 / 0 0.41
Q3 Helps/Neutral/Hinders 16 / 10 / 2 1 / 19 / 2 17 / 29 / 4 0.61
Q4 Yes/No 17 / 11 5 / 17 22 / 28 0.47
Q5 Yes/No 17 / 11 6 / 16 23 / 27 0.80
Table 1: Consolidated label distribution for Q1-Q5. Fleiss’ $\kappa$ scores are computed by averaging each annotator with the consolidated annotation. The consolidated labels and agreement scores are shown for all the samples, as well as partitioned into those where the model prediction was correct/incorrect.
4.3 Module Specific Evaluation
POS Sentiment
Layers BERT RoBERTa BERT RoBERTa
9 92.38 86.97 31.94 99.59
10 92.79 89.64 99.57 99.69
11 93.39 89.95 99.71 99.48
12 93.95 90.04 99.25 99.27
Table 2: Accuracy of PredictionAttributor in mapping a representation to the correct latent concept. See Appendix Table 789 for complete results.
4.3.1 ConceptDiscoverer
ConceptDiscoverer identifies latent concepts by clustering the representation. We questioned whether the discovered latent concepts are a true reflection of the properties that a representation possesses. Using ConceptDiscoverer, we form latent concepts of the last layer and automatically annotate them as described above. We found 87%, 83% and 86% of the latent concepts of BERT, RoBERTa and XLMR that perfectly map to a POS tag respectively. We further analyzed other concepts where 90% of the words did not belong to a single tag. We found them to be of compositional nature i.e. a concept consisting of related semantics like a mix of adjectives and proper nouns about countries such as Swedish and Sweden (Appendix Figure 9). For sentiment, we found 78%, 95% and 94% of the latent concepts of BERT, RoBERTa and XLMR to consist of either Positive or Negative sentences. The high number of class-based clusters of RoBERTa and XLMR show that at 12th layer, majority of their latent space is separated based on these two classes (see Table 6 for detailed results).
Layers 0 2 5 10 12
POS Top 1 100 100 99.03 92.67 84.19
Top 2 100 100 99.75 97.89 94.15
Top 5 100 100 99.94 99.68 99.05
Sentiment Top 1 100 100 97.19 83.09 68.24
Top 2 100 100 99.63 92.67 83.24
Top 5 100 100 99.94 97.75 94.24
4.3.2 PredictionAttributor
We questioned whether the salient input representation correctly represents the latent space of the output. This specifically evaluates PredictionAttributor. We calculate the number of times the representation of the most salient word/[CLS] token maps to the latent concept of the identical label as that of the prediction. We expect a high alignment at the top layers for PredictionAttributor to be correct. We do not include ConceptMapper when evaluating this and conduct the experiment using the training data only where we already know the alignment of a salient representation and the latent concept. Table 2 shows the results across the last four layers (See Appendix Tables 789 for full results). For POS, we observed a successful match of 93.95%, 90.04% and 93.13% for BERT, RoBERTa and XLMR respectively. We observed the mismatched cases and found them to be also of a compositional nature i.e. latent concepts comprised of semantically related words (see Appendix Figure 9 for examples).
4.3.3 ConceptMapper
Here, we evaluate the correctness of ConceptMapper in mapping a test representation to the training data latent concepts. ConceptMapper trains using representations and their cluster ids as labels. For every layer, we randomly split this training data into 90% train and 10% test data. Here, the test data serves as the gold standard annotation of latent concepts. We train ConceptMapper using the training instances and measure the accuracy of the test instances. Table 3 presents the accuracy of the POS and Sentiment tasks using BERT (See Appendix Tables 1011 for results of other models). Observing Top-1 accuracy, the performance of ConceptMapper starts high (100%) for lower layers and drops to 84.19 and 68.24% for the last layer. We found that the latent space becomes dense on the last layer. This is in line with Ethayarajh (2019) who showed that the representations of higher layers are highly anisotropic. This causes several similar concepts close in the space. If true, the correct label should be among the top predictions of the mapper. We empirically tested it by considering the top two and top five predictions of the mapper, achieving a performance of up to 99.05% and 94.24% for the POS and Sentiment tasks respectively.
5 Related work
The explainability methods can be approached by local explanations and global explanations Madsen et al. (2023); Sundararajan et al. (2017a); Denil et al. (2014); Selvaraju et al. (2020); Kapishnikov et al. (2021); Zhao and Aletras (2023); Kim et al. (2018); Ghorbani et al. (2019); Jourdan et al. (2023); Zhao et al. (2023); Ribeiro et al. (2016). Lyu et al. (2023) provides a comprehensive survey on explainability methods in NLP. LACOAT is a local explanation method providing post-hoc explanations given an input instance. One of the most common ways for local explanations is to interpret the model prediction based on the input features. However, the shortcoming of such an explanation is the lack of contextual verbosity, which could not interpret the multifaceted roles of the input features.
A number of works attempted to explain and interpret NLP models using human-defined concepts Kim et al. (2018) and concepts extracted from hidden representations Zhao et al. (2023); Ghorbani et al. (2019); Rajani et al. (2020). Zhao et al. (2023); Kim et al. (2018) worked on the global explanation based on a surrogate model. Zhao et al. (2023) trained the surrogate model using on two optimization criteria i.e. auto encoding loss to stay faithful to the original model distribution and impact of the latent concept to prediction. Different from them, we provide local explanations and we ensure the faithfulness of latent concepts to the model by extracting them directly from the hidden representation without any supervised training. Rajani et al. (2020) used k-nearest neighbors of the training data for low-confidence predictions and showed them to be useful in revealing erroneous correlations and misclassified instances and enhancing the performance of the finetuned model.  Dalvi et al. (2022) analyzed latent concepts of pretrained models in terms of their ability to represent linguistic knowledge. Our ConceptDiscoverer module is motivated from them. However, we propose a method to explain a model’s prediction using training data latent concepts.
6 Conclusion
We presented LACOAT that provides a human-friendly explanation of a model’s prediction. The qualitative evaluation and human evaluation showed that LACOAT explanations are insightful in explaining a correct prediction, in highlighting a wrong prediction and in comparing the explanations of models. The reliance on training data latent space enables interpreting how knowledge is structured in the network. Similarly, it enables the study of the evolution of predictions across layers. LACOAT promises human-in-the-loop in the decision-making process and is a step towards trust in AI.
7 Limitations
A few limitations of LACOAT are: 1) while hierarchical clustering is better than nearest neighbor in discovering latent concepts as established by Dalvi et al. (2022), it has computational limitations and it can not be easily extended to a corpus of say 1M tokens. However, the assumptions that are taken in the experimental setup e.g. considering the maximum 20 occurrences of a word (supported by Dalvi et al. (2022)) work well in practice in terms of limiting the number of tokens and covering all facets of a majority of the words. Moreover, the majority of the real-world tasks have limited task-specific data and LACOAT can effectively be applied in such cases.
2) LACOAT requires white-box access to the training data and the model for explanation which may not be available for large language models such as Llama and ChatGPT. For open-source models like Llama and Mistral, LACOAT can be applied in their task-specific usage. For instance for sentiment classification using Llama, the training data of sentiment can be used to extract the latent space of Llama for the sentiment classification task. However, the applicability of LACOAT for generative tasks is limited due to the open vocabulary.
Appendix A Salient Representation Extraction - Position Based Attribution Method
This strategy uses the position of the output head as an indication of the most salient contextual representation. For instance,
  • In the case of sequence classification, the representation of the [CLS] token, $\overrightarrow{z}_{\texttt{[CLS]}}$ (or a model’s representative classification token) will be considered as the most salient representation.
  • In the case of masked token prediction, the representation of the [MASK] token ($\overrightarrow{z}_{\texttt{MASK}}$) will be considered as the most salient for making the prediction.
  • In the case of sequence labeling, the representation at the time step of the prediction will be used. For example, in the case of POS tagging, for the prediction of a tag of the word love in the sentence I [love] soccer, the second time step’s representations ($\overrightarrow{z}_{w_{2}}$) will be used.
Appendix B Finetuning Performance
We tuned several transformers BERT-base-cased, RoBERTa and XLM-RoBERTa. We used standard splits for training, development and test data that we used to carry out our analysis. The splits to preprocess the data are available through git repository [^1] https://github.com/nelson-liu/contextual-repr-analysis. See Table 4 and Table 5 for statistics and classifier accuracy.
@@ Appendix B Finetuning Performance
Table 4: The fine-tuned performance of models, data statistics (number of sentences) on training, development, and test sets used in the experiments, and the number of tags to be predicted for the POS tagging task. Model: BERT, RoBERTa, XLM-R
@@
Table 5: The fine-tuned performance of models, data statistics (number of sentences) on training, development, and test sets used in the experiments, and the number of tags to be predicted for the sentiment classification task. Model: BERT, RoBERTa, XLM-R
Appendix C Qualitative Evaluation - More Examples
C.1 Example for the Evolution of Concepts
5 presents the other example of latent concepts of the salient words in layer 0, 6, and 12. Similarly to the example shown in 2, the latent concept of this example shows that the different forms of the verb “sit” and are not aligned with its usage in the input instance. The concept in the middle layer aligns better with the sentiment of the input sentence(Figure 5(b)). The most words of layer 6’s latent concept match the sentiment of the input sentence. We also randomly pick five [CLS] instances from the latent concept and show their corresponding sentences in the figure (see Figure 5(c)). The concept of the last layer is best aligned with the input sentence.
@@ C.1 Example for the Evolution of Concepts
C.2 Adversarial Example of the Sentence in Figure 3(b)
@@
C.3 Correct Predicted Label with Incorrect Gold Label
@@
C.4 Incorrect Prediction in POS tagging Task
@@
Appendix D Module Specific Evaluation
D.1 ConceptDiscoverer - Compositional Concept Examples
@@
D.2 ConceptDiscoverer - Number of Clusters For Each Polarity in the Sentiment Classification Task
@@
D.3 ConceptMapper - Position Based Attribution Method
We evaluate the module that is based on the position based attribution approach as the number of times an input representation at the position of the output head maps to the latent concept that is annotated with an identical label as the output. For example, consider that the model predicts Proper Noun (PN) for the input word "Trump". In order for the input representation of the predicted label to be aligned with the latent concept, the representation of the word "Trump" on at least the last layer should be in a cluster of words whose label is PN. [^2] We labeled concepts with a tag if 90% of the words in the concept belongs to one class. Similarly for sentiment classification, we expect the [CLS] representation on the last layer to map to a latent concept that is dominated by the same class as the prediction.
For POS tagging task, the salient representation is identical for both the position based and saliency based methods and results in the same performance (Table 7). For the sentiment classification task, performance drops and even reaches zero as in the case of XLMR. This may be due to the position-based method which fails to find the right latent concept when the most attributed word is different from the position of the output head (Table 8).
@@ D.3 ConceptMapper - Position Based Attribution Method
Layer 0 16.81 14.29 17.66
Layer 1 17.79 16.49 18.89
Layer 2 21.16 20.18 20.71
Layer 3 22.79 20.13 31.03
Layer 4 29.70 24.65 40.51
Layer 5 46.74 29.26 60.31
Layer 6 73.19 42.38 77.32
Layer 7 84.52 57.46 85.78
Layer 8 90.68 82.84 89.41
Layer 9 92.38 86.97 91.97
Layer 10 92.79 89.64 92.64
Layer 11 93.39 89.95 92.59
Layer 12 93.95 90.04 93.13
Table 7: Position/Saliency-based method: accuracy of PredictionAttributor in mapping a representation to the correct latent concept in the POS tagging task.
Sentiment
Layer BERT RoBERTa XLM-R
Layer 0 0 0 0
Layer 1 0 0 0
Layer 2 0 0 0
Layer 3 0 0 0
Layer 4 0 0 0
Layer 5 0 0 0
Layer 6 0 0 0
Layer 7 0 0 0
Layer 8 0 99.11 0
Layer 9 37.09 98.45 0
Layer 10 99.55 99.14 0
Layer 11 99.82 99.27 99.17
Layer 12 99.25 99.27 99.08
Table 8: Position-based method: accuracy of PredictionAttributor in mapping a representation to the correct latent concept in the Sentiment Classification task. The reason of zero values is that the position-based method fails to find the right latent concept when the most attributed word is different from the position of the output head.
@@
D.4 ConceptMapper - Accuracy of ConceptMapper for the Sentiment Classification and POS Tagging task

### v2

@@ Appendix A Task-specific Prompts used with PlausiFyer
Appendix B Datasets
Task Train Dev Tags
Sentiment 13878 856 2
POS 36557 1802 48
Toxicity 9000 800 2
MNLI 9000 1200 3
Table 4: The data statistics of each dataset used in the evaluation experiments and the number of tags to be predicted. POS Marcus et al. (1993), Jigsaw Toxicity dataset cjadams et al. (2017), the ERASER Sentiment dataset Pang and Lee (2004); Zaidan and Eisner (2008) and the MNLI dataset Wang et al. (2019)
Appendix C Finetuning Performance
We tuned several transformers BERT-base-cased, RoBERTa and XLM-RoBERTa. We used standard splits for training, development and test data that we used to carry out our analysis. The splits to preprocess the data are available through git repository [^2] https://github.com/nelson-liu/contextual-repr-analysis. See Table 5 and Table 6 for statistics and classifier accuracy. We present the results of Toxicity and MNLI in Appendix H and I.
@@ Appendix C Finetuning Performance
Table 5: The fine-tuned performance of models, data statistics (number of sentences) on training, development, and test sets used in the finetuning, and the number of tags to be predicted for the POS tagging task. Model: BERT, RoBERTa, XLM-R
@@
Table 6: The fine-tuned performance of models, data statistics (number of sentences) on training, development, and test sets used in the finetuning, and the number of tags to be predicted for the sentiment classification task. Model: BERT, RoBERTa, XLM-R
Appendix D Qualitative Evaluation - More Examples
D.1 Example for the Evolution of Concepts
Figure 5 presents the other example of latent concepts of the salient words in layers 0, 6, and 12. Similarly to the example shown in Figure 2, the latent concept of this example shows that the different forms of the verb “sit” are not aligned with its usage in the input instance. The concept in the middle layer aligns better with the sentiment of the input sentence (Figure 5(b)). Most words of layer 6’s latent concept match the sentiment of the input sentence. We also randomly pick five [CLS] instances from the latent concept and show their corresponding sentences in the figure (see Figure 5(c)). The concept of the last layer is best aligned with the input sentence.
@@ D.1 Example for the Evolution of Concepts
D.2 Adversarial Example of the Sentence in Figure 3(b)
@@
D.3 Correct Predicted Label with Incorrect Gold Label
@@
D.4 Incorrect Prediction in POS tagging Task
@@
Appendix E Module Specific Evaluation
E.1 ConceptDiscoverer - Compositional Concept Examples
@@
E.2 ConceptDiscoverer - Number of Clusters For Each Polarity in the Sentiment Classification Task
@@
Layer 0 13.76 11.13 11.97
Layer 1 12.75 13.58 11.91
Layer 2 15.51 15.60 12.99
Layer 3 17.61 17.25 22.88
Layer 4 23.81 20.30 32.08
Layer 5 37.03 23.28 48.44
Layer 6 64.83 32.52 67.94
Layer 7 77.90 48.61 80.11
Layer 8 86.96 73.88 85.83
Layer 9 88.98 82.56 89.30
Layer 10 89.99 83.24 89.94
Layer 11 90.68 84.61 90.19
Layer 12 92.16 85.67 90.18
Table 8: Saliency-based method (95%): accuracy of PredictionAttributor in mapping a representation to the correct latent concept in the POS tagging task. Model: BERT-base-cased, RoBERT-base, XLM-R
@@
E.3 ConceptMapper - Accuracy of ConceptMapper for the Sentiment Classification and POS Tagging task